a marketing manager makes the statement that the long-run probability that a customer will the deluxe model to the standard model of product is 30%.what is the probability that more than tow in random sample of 10 customers will prefer the standard model?
1
Expert's answer
2012-08-30T10:02:29-0400

Probability is 30%*2/10=6%
There is probability that two out of 10 prefer standard model. It`s probability = 2/10 = 0.2
Thus to calculate total probability we need to multiply 0.2 and 30%